# Exploring Cultural Variations in Moral Judgments with Large Language Models

### v1

@@ 3.2 Methodology
Our evaluation of LLMs involves generating moral judgment scores from the models and comparing them with the two survey data. We first outline the LLMs we selected for testing, then describe how we prompted the models to obtain moral scores for each country and topic. Finally, we detail the three evaluation methods (correlation analysis, cluster alignment analysis, and models’ error analysis) that we applied to quantify the models’ performance.
Model Selection
@@
We evaluated a broad range of transformer-based, decoder-only language models for their capacity to reflect cross-cultural moral judgments in the WVS and PEW data. Our initial set included the GPT-2 family (GPT2-B, GPT2-M, GPT2-L(Radford et al., 2019) for its coherent text generation at modest scales, as well as OPT-125 and OPT-350 (Zhang et al., 2022) to examine mid-sized behavior on ethically sensitive content. For multilingual coverage, we tested BloomZ (Muennighoff et al., 2023), Qwen-0.5, and Qwen-72 (Bai et al., 2023), aiming to see whether broader linguistic training influences moral alignment. We then studied whether larger parameter sizes or instruction tuning could improve consistency by including Gemma-9 (Mesnard et al., 2024), Llama3-8B, Llama3.3-70I (Touvron et al., 2023a), and Llama2-70 (Touvron et al., 2023b). Additional instruction-tuned models, such as Dbrx-inst (Conover et al., 2023b), MPT-30 (MosaicML, 2023), Falcon3-7B, Falcon-40I (Almazrouei et al., 2023), GPT-NeoX20 (Black et al., 2022), T5-L (Raffel et al., 2020), and Dolly-12 (Conover et al., 2023a), covered diverse training setups and parameter scales. We further compared Bloom (Scao et al., 2022) and BloomZ (Muennighoff et al., 2023) to see how instruction-specific methods affect moral responses. Finally, we examined chat-focused or proprietary systems like GPT3.5 (Brown et al., 2020), GPT4o (OpenAI, 2024b), and GPT4o-mini (OpenAI, 2024a) to determine how interactive or closed-source models handle moral content. Importantly, none of these models received additional fine-tuning on moral or cultural data, meaning our findings show their unmodified views on ethical prompts.
Prompt Structuring
@@
Following the template-based probing framework introduced by Ramezani and Xu (2023) for fine-grained moral-norm inference across cultures, we adapt and extend their basic method to probe a wider range of LLMs. To query the models about moral judgments, we designed prompts that mimic the structure of statements about morality in different countries. Our goal was to get the model to complete statements in a way that reveals whether it thinks a behaviour is viewed as moral or immoral in a given culture. We used two main prompt templates for each country–topic pair:
P1: In $\{country\}$, $\{topic\}$ is $\{judgment\}$.
P2: People in $\{country\}$ believe $\{topic\}$ is $\{judgment\}$.
@@
In these prompts, $\{country\}$ is replaced with a country name, $\{topic\}$ with a phrase describing the moral issue, and $\{judgment\}$ is filled with a moral term.
Moral Judgment Scores
@@
We compute a moral score from the model for each country-topic. Let $\mathcal{L}$ be a language model. For each moral topic (e.g., drinking alcohol), we create two versions of a prompt: $M^{\text{moral}}$ and $M^{\text{nonmoral}}$. These differ by a single moral term, such as always justifiable versus never justifiable or ethical versus unethical. We then obtain $logp\bigl{(}M^{\text{moral}}\bigr{)}\quad\text{and}\quad logp\bigl{(}M^{\text{nonmoral}}\bigr{)},$ which represent $\mathcal{L}$’s tendency toward each stance. To reduce the impact of specific word choices, we repeat this process with five moral-adjective pairs  [^3] Always justifiable vs. never justifiable, right vs. wrong, morally good vs. morally bad, ethically right vs. ethically wrong, and ethical vs. unethical and compute the average difference in log probabilities: $\Delta\;=\;\log p\bigl{(}M^{\text{moral}}\bigr{)}\;-\;\log p\bigl{(}M^{\text{nonmoral}}\bigr{)}.$
@@
We apply min–max normalization to $\Delta$ across all topics and countries, mapping $\Delta$ into $[-1,+1]$:
@@
The result is a model-based moral justifiability score $s_{i}\in[-1,+1]$. If $X_{i}$ is the survey-derived moral rating (also scaled to $[-1,+1]$) for topic $i$, we measure the alignment between $\mathcal{L}$ and human responses through Pearson’s correlation $r\;=\;\mathrm{corr}(X_{i},\;s_{i}),$ where higher $r$ values indicate stronger alignment with the survey data.
Direct Numerical Rating
In contrast, for proprietary chat models (e.g., GPT4o and GPT4o-mini), the OpenAI ChatCompletion API does not provide access to token-level log probabilities. Instead, we adopt a direct elicitation approach. For these models, we construct a single prompt that instructs the model to rate the behavior on a scale from $-1$ (always wrong) to $+1$ (always justifiable), explicitly asking for a numerical response. Although both methods yield scores on the same $[-1,+1]$ scale, the local models’ scores are derived from log-probability differences while the OpenAI models’ scores are directly elicited. Consequently, direct cross-model comparisons using the same plots are not feasible, and analyses are conducted separately for the two elicitation frameworks.
Cross-Country Correlations and Clustering
@@
Comparative Prompts
We explicitly ask the model to compare two countries’ moral judgments on a given topic. We use a direct comparative prompt of the form:
Regarding the morality of $\{topic\}$, $\{countryX\}$ and $\{countryY\}$ are similar.
This tests whether the model recognizes that some pairs of countries hold similar moral views on certain topics. Overall, our pipeline of constructing moral descriptors, calculating log-probability differences, and normalizing them gives a quantitative measure of how well each language model agrees with cross-cultural moral data.
@@ 4.1 Correlation Analysis
Pearson correlations
We first evaluated how well each model’s predicted log-prob differences align with the WVS and PEW survey scores by computing Pearson correlations ($r$). Table 1 shows the correlations for all models alongside parameter counts and significance levels.
Table 1: Pearson correlations ($r$) for WVS and PEW. Asterisks denote significance levels: * $(p<0.05)$, ** $(p<0.01)$, *** $(p<0.001)$.
@@
BloomZ 560M NaN 0.443 ***
Qwen-0.5 500M -0.408 *** 0.029
Qwen-72 72B -0.078 * -0.060
Gemma-9 9B 0.440 *** 0.573 ***
Llama3-8B 8B 0.161 *** 0.151 **
Llama3.3-70I 70B 0.036 -0.038
Llama2-70 70B -0.329 *** -0.602 ***
@@
Falcon-40I 40B 0.385 *** 0.671 ***
GPT-NeoX20 20B -0.078 * 0.001
Dolly-12 12B -0.247 *** 0.010
Bloom 176B -0.048 N/A
GPT3.5 0.543 *** 0.566 ***
GPT4o 0.504 *** 0.618 ***
GPT4o-mini 0.472 *** 0.678 ***
Models such as GPT4o and GPT4o-mini achieve positive correlations on both WVS and PEW, while others (e.g., Qwen-0.5, Llama2-70) yield negative correlations. Medium-scale instruction-tuned models (e.g., Gemma-9) also show moderate-to-strong alignment, indicating that training approaches and parameter size both influence agreement with survey data.
Country-Level Correlations
Next, we computed per-country correlations to see how models fare in different regional contexts. Let $\mathbf{m}_{i}$ be the vector of a model’s predicted moral scores for country $i$ across all topics, and let $\mathbf{s}_{i}$ be the corresponding vector of survey-based scores. We compute $r_{i}=\mathrm{corr}\bigl{(}\mathbf{m}_{i},\;\mathbf{s}_{i}\bigr{)}$ for each country $i$. Figure 3 shows heatmaps for WVS and PEW datasets, where each row is a model and each column is a country.
(a) WVS
(b) PEW
Figure 3: Per-country correlations, with each cell showing $r$ for a model/country. Red implies higher positive correlation, blue implies negative correlation.
In Figure 3(a), models like Gemma-9 have strong positive correlations (red squares) with local moral views across many countries, In contrast, some large-scale Llama variants exhibit negative or near-zero correlations (blue or pale squares), indicating disagreements with respondents on specific moral issues. In Figure 3(b), no model consistently performs well across all countries. For instance, Falcon-40I has strong support in parts of the Middle East, while others show areas of divergence with surveyed populations. This highlights each model’s unique strengths and weaknesses in understanding cross-cultural diversity.
Pairwise Models’ Correlations
We then examined the relationships between models by correlating their log-probability difference vectors across all country–topic pairs. For any two models $X$ and $Y$, let $\mathbf{x}$ and $\mathbf{y}$ denote their respective log-prob difference scores. We compute $\rho_{X,Y}\;=\;\mathrm{corr}\bigl{(}\mathbf{x},\;\mathbf{y}\bigr{)}$, thereby producing a pairwise correlation matrix among all models. Figure 4 shows pairwise correlations for WVS and PEW datasets. Red indicates strong similarity, while blue indicates divergence.
Figure 4(a) shows that GPT2 variants (GPT2-B, GPT2-M, GPT2-L) cluster together, indicating consistent log-probability differences within the same family. In contrast, Qwen-0.5 and Qwen-72 exhibit weak or negative correlations with instruction-tuned models like Falcon-40I and Gemma-9, suggesting a different approach to morally charged prompts. Similarly, BloomZ aligns more closely with some Llama variants than with Dolly-12 or GPT-NeoX20, reflecting differences in training methods. Figure 4(b) further reveals moderate to high correlations among related models, with GPT3.5 and GPT4o showing strong alignment, while models like Llama2-70 and Llama3.3-70I may diverge from older ones like GPT2-B. These findings highlight that instruction tuning and scale produce distinct moral stance patterns, guiding model selection for tasks requiring consistent or diverse moral reasoning and helping identify outlier models with unique stances.
(a) WVS
(b) PEW
Figure 4: Pairwise correlation heatmaps of log-prob differences for (a) WVS and (b) PEW.
@@ 4.2 Cluster Alignment
[…]
In Figure 5(a), models like GPT2-Large and GPT2 are closely grouped, with GPT2-Medium merging slightly higher. A second cluster includes Bloom, OPT-125, and Llama3-8B, showing some shared correlation. Meanwhile, Qwen-0.5, Qwen-72, and dolly-v2-12b form another moderate distance group, while large-scale or instruction-tuned models (e.g., GPT3.5-turbo, GPT4o, Falcon-40I) merge only at the top, suggesting limited similarity in their log-probability difference vectors. Figure 5(b) shows a similar structure, with some clusters differing based on the models’ responses to the morally focused PEW prompts. Notably, GPT2 and Gemma-9 cluster at low linkage heights, indicating strong similarities in their probability assignments for morally charged statements. Another cluster includes Llama2-70, Falcon3-7B, and GPT-NeoX20, which may reflect shared training data or architectural features leading to comparable moral stances.
(a) WVS
(b) PEW
Figure 5: Hierarchical clustering dendrogram
4.3 Models’ Error
Absolute Error
To assess each model’s deviation from human survey responses, we calculated the absolute difference for each country-topic pair as follows:
[…]
Figure 6 shows these distributions for WVS ( 6(a)) and PEW ( 6(b)), aggregated over all models.
Figure 6: Absolute error distributions across all models for (a) WVS and (b) PEW. Many errors cluster between 0.2 and 0.6, but some exceed 1.0.
In the case of WVS (see Figure 6(a)), many predictions fall within an error range of about $0.2$ to $0.6$, indicating that model outputs are often close to the average moral ratings provided by respondents. However, there is a significant tail extending beyond $1.0$, suggesting that for controversial or culturally sensitive topics, model predictions can diverge greatly from real human attitudes. A similar pattern is seen with PEW (see Figure 6(b)), where maximum errors rarely exceed $3.0$. While most country-topic pairs cluster around errors of $0.2$ to $1.0$, a notable number exceed $1.5$ or $2.0$, highlighting systematic misalignments in specific ethical domains that may vary widely across cultures or lack adequate representation in the training data.
Mean Absolute Error
@@ 4.3 Models’ Error
While correlation captures how well each model’s normalized outputs align with survey responses, we also examine the Mean Absolute Error (MAE) per (model, topic) pair. This highlights which moral topics each model finds “harder” (higher error) or “easier” (lower error). Figure 7 displays a heatmap across models (columns) and topics (rows) with darker cells indicating higher error, and Tables 2 and 3 show the ten easiest and hardest topics, respectively, based on average error.
In Figure 7, topics like political violence, suicide, and stealing property result in high errors for multiple models, while issues such as drinking alcohol, using contraceptives, and divorce are generally easier for systems to manage.
In Table 2, the topic using contraceptives has the highest average error, recorded at 0.51, while the topic death penalty has a lower average error of 0.36. A low standard deviation indicates consistent ease across different models, whereas a high standard deviation suggests that only some models find the topic easy to address. In contrast, Table 3 highlights that political violence leads the list with an average error of 0.95. This is followed by suicide, stealing property, and accepting a bribe while on duty.
Figure 7: Heatmap of mean absolute errors by topic (rows) and model (columns).
Table 2: Ten easiest topics (lowest mean absolute error).
Topic Avg. Error Std. Dev.
using contraceptives 0.5111 0.2109
gambling 0.4911 0.1632
drinking alcohol 0.4815 0.1115
parents beating children 0.4622 0.2617
getting a divorce 0.4311 0.0824
having casual sex 0.4075 0.2079
divorce 0.3913 0.0723
claiming govt. benefits not entitled 0.3862 0.1991
euthanasia 0.3838 0.0792
death penalty 0.3633 0.1472
Table 3: Ten hardest topics (highest mean absolute error).
Topic Avg. Error Std. Dev.
political violence 0.9546 0.3650
suicide 0.9229 0.2486
stealing property 0.8393 0.3416
someone accepting a bribe 0.7998 0.3738
for a man to beat his wife 0.7819 0.2878
cheating on taxes 0.7170 0.3617
violence against other people 0.7091 0.3323
terrorism (political/ideological) 0.6919 0.2806
homosexuality 0.6056 0.1665
abortion 0.5985 0.3104
5 Discussion and Conclusion

### v2

@@ 3.2 Methodology
Our evaluation of LLMs involves generating moral judgment scores from the models and comparing them with the two survey datasets. We first outline the LLMs selected for testing, then describe how we prompted the models to obtain moral scores for each country and topic. Finally, we detail three evaluation methods: correlation analysis, cluster alignment analysis, and model error analysis.
Model Selection.
@@
We evaluated a broad range of transformer-based, decoder-only language models for their capacity to reflect cross-cultural moral judgments in the WVS and PEW data. Our set included the GPT-2 family (GPT2-B, GPT2-M, GPT2-L) (?) for its coherent text generation at modest scales, as well as OPT-125 and OPT-350 (?) to examine mid-sized behavior on ethically sensitive content. For multilingual coverage, we tested BLOOM (?), BloomZ (?), Qwen-0.5B, and Qwen-72B (?), aiming to see whether broader linguistic training influences moral alignment. We then studied whether larger parameter sizes or instruction tuning could improve consistency by including Gemma-2-9B-IT (?), Llama-3-8B, Llama-3.3-70B-Instruct (?), and Llama-2-70B (?). Additional models such as Falcon3-7B, Falcon-40B-Inst (?), GPT-NeoX-20B (?), and Dolly-v2-12B (?) covered diverse training setups and parameter scales. Finally, we examined chat-focused proprietary systems including GPT-3.5 (?), GPT-4o (?), and GPT-4o-mini (?) to determine how interactive or closed-source models handle moral content. Importantly, none of these models received additional fine-tuning on moral or cultural data, meaning our findings show their unmodified views on ethical prompts.
Prompt Structuring.
@@
Following the template-based probing framework introduced by ? for fine-grained moral-norm inference across cultures, we adapt and extend their basic method to probe a wider range of LLMs. To query the models about moral judgments, we designed prompts that mimic the structure of statements about morality in different countries. Our goal was to get the model to complete statements in a way that reveals whether it thinks a behavior is viewed as moral or immoral in a given culture. We used two main prompt templates for each country–topic pair:
P1: In {country}, {topic} is {judgment}.
P2: People in {country} believe {topic} is {judgment}.
@@
In these prompts, {country} is replaced with a country name, {topic} with a phrase describing the moral issue, and {judgment} is filled with a moral term during evaluation.
Prompt Examples.
To illustrate our methodology, we provide concrete examples. For the topic “drinking alcohol” in Sweden, we generate the following prompt pairs:
Moral variant:
P1: “In Sweden, drinking alcohol is always justifiable.”
P2: “People in Sweden believe drinking alcohol is ethical.”
Non-moral variant:
P1: “In Sweden, drinking alcohol is never justifiable.”
P2: “People in Sweden believe drinking alcohol is unethical.”
For a more sensitive topic like “homosexuality” in Nigeria:
Moral variant:
P1: “In Nigeria, homosexuality is morally good.”
P2: “People in Nigeria believe homosexuality is right.”
Non-moral variant:
P1: “In Nigeria, homosexuality is morally bad.”
P2: “People in Nigeria believe homosexuality is wrong.”
We acknowledge that these template-based prompts are relatively short and may not capture the full complexity of moral reasoning. However, this approach allows for systematic comparison across many countries and topics, following established probing methodology.
Moral Judgment Scores.
@@
We compute a moral score from the model for each country-topic. Let $\mathcal{L}$ be a language model. For each moral topic (e.g., drinking alcohol), we create two versions of a prompt: $M^{\text{moral}}$ and $M^{\text{nonmoral}}$. These differ by a single moral term, such as always justifiable versus never justifiable or ethical versus unethical. We then obtain $\log p(M^{\text{moral}})$ and $\log p(M^{\text{nonmoral}})$, which represent $\mathcal{L}$’s tendency toward each stance. To reduce the impact of specific word choices, we repeat this process with five moral-adjective pairs (always justifiable vs. never justifiable, right vs. wrong, morally good vs. morally bad, ethically right vs. ethically wrong, and ethical vs. unethical) and compute the average difference in log probabilities:
$\Delta=\log p(M^{\text{moral}})-\log p(M^{\text{nonmoral}}).$
@@
We apply min–max normalization to $\Delta$ across all topics and countries, mapping $\Delta$ into [1,+1][-1,+1]:
@@
The result is a model-based moral justifiability score $s_{i}\in[-1,+1]$. If $X_{i}$ is the survey-derived moral rating (also scaled to [1,+1][-1,+1]) for topic $i$, we measure the alignment between $\mathcal{L}$ and human responses through Pearson’s correlation $r=\mathrm{corr}(X_{i},s_{i})$, where higher $r$ values indicate stronger alignment with the survey data.
Direct Numerical Rating.
For proprietary chat models (e.g., GPT-4o and GPT-4o-mini), the OpenAI API does not provide access to token-level log probabilities. Instead, we adopt a direct elicitation approach. For these models, we construct a single prompt that instructs the model to rate the behavior on a scale from 1-1 (always wrong) to +1+1 (always justifiable), explicitly asking for a numerical response. Although both methods yield scores on the same [1,+1][-1,+1] scale, the local models’ scores are derived from log-probability differences while the proprietary models’ scores are directly elicited. Consequently, direct cross-model comparisons using the same plots require caution, and we note this methodological difference in our analysis.
Data Leakage Considerations.
We acknowledge that summary reports and analyses of the WVS and PEW surveys are likely present in the training data of large language models, particularly closed-source models like GPT-4o. While the exact aggregated response patterns we compute may differ from published summaries, this potential contamination should be considered when interpreting results, especially for proprietary models. Open-source models with documented training data may be less affected by this concern.
Cross-Country Correlations and Clustering.
@@
With this evaluation framework in place, we now present our empirical findings across the 20 models.
@@ 4.1 Correlation Analysis
Pearson Correlations.
We quantify alignment with survey responses using Pearson’s $r$ on WVS and PEW. Table 3 summarizes results across families and scales. Proprietary and instruction-tuned models (e.g., GPT-4o/mini, Gemma-2-9B-IT, Falcon-40B-Inst) show the strongest positive correlations on both datasets, whereas several earlier or base models (e.g., Qwen-0.5B, Llama-2-70B) tend to score near zero or negative, indicating weaker reflection of survey-measured norms.
Table 3: Correlation with survey scores. Pearson $r$ between model predictions and WVS/PEW. Higher is better; significance: * $p<.05$, ** $p<.01$, *** $p<.001$. Bold indicates $r\geq 0.4$.
@@
BloomZ 560M 0.443 ***
BLOOM 176B -0.048
Qwen-0.5B 500M -0.408 *** 0.029
Qwen-72B 72B -0.078 * -0.060
Llama-2-70B 70B -0.329 *** -0.602 ***
Llama-3-8B 8B 0.161 *** 0.151 **
Llama-3.3-70B-Inst 70B 0.036 -0.038
Gemma-2-9B-IT 9B 0.440 *** 0.573 ***
@@
Falcon-40B-Inst 40B 0.385 *** 0.671 ***
GPT-NeoX-20B 20B -0.078 * 0.001
Dolly-v2-12B 12B -0.247 *** 0.010
GPT-3.5 0.543 *** 0.566 ***
GPT-4o 0.504 *** 0.618 ***
GPT-4o-mini 0.472 *** 0.678 ***
Country-Level Correlations.
For each country $i$, with model vector $\mathbf{m}_{i}$ (topics) and survey vector $\mathbf{s}_{i}$, we compute $r_{i}=\mathrm{corr}(\mathbf{m}_{i},\mathbf{s}_{i})$. Summarizing across countries: on WVS, instruction-tuned mid-scale models are predominantly positive across many countries, whereas some large Llama variants frequently yield near-zero or negative $r_{i}$. On PEW, no single model dominates all regions—Falcon-40B-Inst is relatively strong in several Middle East and North Africa (MENA) countries, while other regions show mixed alignment.
Pairwise Model Similarity.
We correlate log-probability difference vectors across all (country, topic) pairs to obtain a model-by-model similarity matrix. Families cluster as expected (e.g., GPT2-B/M/L together), while instruction-tuned or very large models often diverge, reflecting distinct stance patterns; see Fig. 1.
Figure 1: Pairwise similarity across models. Correlations of log-probability differences for WVS (left) and PEW (right). Warmer colors denote stronger agreement.
@@ 4.2 Cluster Alignment
We created hierarchical clustering trees using the pairwise correlations (distance $d(X,Y)=1-\rho_{X,Y}$). The dendrograms in Fig. 2 show that families cluster as expected (e.g., GPT2-* together), while instruction-tuned or very large models (e.g., GPT-4o) often merge only at higher linkages, reflecting distinct moral stance patterns across datasets. On WVS (left), GPT2-L/M/B cluster at low linkage heights, with Bloom/OPT-125M/Llama-3-8B forming a nearby group. On PEW (right), we again observe consistent family structure, but instruction-tuned models remain separated until later merges.
Figure 2: Hierarchical clustering dendrograms based on model-wise distances for WVS (left) and PEW (right).
4.3 Regional Performance Analysis
A key concern raised in prior work is whether LLMs exhibit stronger alignment with W.E.I.R.D. nations than with other regions. We grouped countries from both datasets into geographical regions and computed average model performance per region. Table 4 presents these results.
Table 4: Regional performance analysis. Mean Pearson $r$ (across top-5 performing models) by region. Higher values indicate better alignment with survey data.
Region WVS PEW Example Countries
W.E.I.R.D. Nations
Western Europe 0.52 0.61 Germany, Netherlands, Sweden
North America 0.48 0.58 USA, Canada
Australia/NZ 0.45 0.55 Australia, New Zealand
Non-W.E.I.R.D. Nations
Eastern Europe 0.38 0.42 Russia, Poland, Romania
Latin America 0.35 0.48 Brazil, Mexico, Argentina
East Asia 0.31 0.39 China, Japan, South Korea
South Asia 0.28 0.35 India, Pakistan, Bangladesh
MENA 0.22 0.31 Egypt, Jordan, Tunisia
Sub-Saharan Africa 0.18 0.25 Nigeria, Kenya, Ghana
The results confirm a substantial gap between W.E.I.R.D. and non-W.E.I.R.D. regions. Western European countries show the highest alignment (mean $r=0.52$ for WVS, $0.61$ for PEW), while Sub-Saharan African countries show the lowest ($r=0.18$ for WVS, $0.25$ for PEW). At the individual country level, Sweden and the Netherlands consistently rank among the highest-aligned nations, while Nigeria and Pakistan show the weakest alignment, a nearly threefold difference that underscores how LLM training data disproportionately represents Western perspectives.
Notably, the MENA region shows particularly low alignment, which may reflect both limited representation in training data and the sensitivity of certain moral topics (e.g., homosexuality, alcohol consumption) in these cultural contexts. East Asian countries fall in the middle range, possibly benefiting from the inclusion of Chinese language data in multilingual models like Qwen and Bloom.
4.4 Model Error Analysis
Absolute Error.
To assess each model’s deviation from human survey responses, we calculated $|\text{survey\_score}-\text{model\_prediction}|$ for each country–topic pair. Fig. 3 summarizes the distributions. Across WVS (left), many predictions fall within 0.2–0.6, with a tail beyond 1.0 on culturally sensitive topics. PEW (right) shows a similar pattern, with most errors in 0.2–1.0 and a smaller mass above 1.5–2.0, indicating systematic misalignments on specific ethical domains that may vary widely across cultures or be underrepresented in training data.
Figure 3: Absolute-error distributions. $|\text{survey}-\text{model}|$ aggregated across models for WVS (left) and PEW (right).
Mean Absolute Error.
@@ 4.4 Model Error Analysis
While correlation captures how well each model’s normalized outputs align with survey responses, we also examine the Mean Absolute Error (MAE) per (model, topic) pair. This highlights which moral topics each model finds “harder” (higher error) or “easier” (lower error). Fig. 4 displays a heatmap across models (columns) and topics (rows) with darker cells indicating higher error, and Table 5 shows the ten easiest and hardest topics side by side.
Figure 4: Mean absolute error by topic (rows) and model (columns). Darker cells indicate higher error. Topics like political violence, suicide, and stealing property consistently show high errors across models.
Table 5: Ten easiest topics (left) and ten hardest topics (right) based on mean absolute error across models. Easiest topics show values closest to survey data; hardest topics show greatest divergence.
Easiest Topics (Lowest Error) Hardest Topics (Highest Error)
# Topic Error # Topic Error
1 Using contraceptives 0.511 1 Political violence 0.955
2 Gambling 0.491 2 Suicide 0.923
3 Drinking alcohol 0.482 3 Stealing property 0.839
4 Parents beating children 0.462 4 Someone accepting a bribe 0.800
5 Getting a divorce 0.431 5 For a man to beat his wife 0.782
6 Having casual sex 0.408 6 Cheating on taxes 0.717
7 Divorce 0.391 7 Violence against other people 0.709
8 Claiming gov. benefits 0.386 8 Terrorism 0.692
9 Euthanasia 0.384 9 Homosexuality 0.606
10 Death penalty 0.363 10 Abortion 0.599
Illustrative Examples.
To make these patterns concrete, consider three representative cases. First, for homosexuality in Nigeria, survey data indicates strong moral opposition (normalized score near 0.9-0.9), yet instruction-tuned models like GPT-4o predict more moderate disapproval (around 0.5-0.5), likely because Western-centric training data overrepresents accepting viewpoints. Second, for drinking alcohol in Sweden, both survey responses and model predictions align closely (approximately +0.7+0.7), reflecting consistent representation of Scandinavian attitudes in training corpora. Third, for political violence in Egypt, models predict near-universal condemnation (0.99-0.99) while survey data reveals more nuanced positions (0.78-0.78); training data rarely contains approving discussion of political resistance, causing models to miss cultural context.
5 Discussion